# The mantle-inner core gravitational mode of oscillation in a strong magnetic field regime

### v1

@@ 4 Results
Figure 2: The angular velocity of the inner core $\Omega_{i}$ (thick orange) and of the fluid core $\Omega_{f}$ (blue) as a function of cylindrical radius ($s/r_{f}$) for the modes identified by a purple outline on the respective panels of Figure 1. In each panel, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
Since we expect the $B_{s}$-field strength in the core to be of the order of a few mT, Alfvén waves should propagate through the ROTC in a short timescale of a couple of years, more rapidly than their timescale of attenuation by EM coupling at the CMB (which should be of the order of a few decades e.g. dumberry08c; gillet10). The MICG mode should then not exist as a distinct mode of oscillation; rather, we expect it to be absorbed into the spectrum of TO modes, although the TO modes are influenced by gravitational coupling.
To further explore how gravitational coupling affects TO modes, let us restore $G_{m}=10^{8}$ S and now set $\overline{B}_{s}=3$ mT. With these choices, the MICG mode is not an independent mode and the first TO harmonic has a period of approximately 6 years. Figure 3 shows how the period (panel a) and $Q$ (panel b) of the first four TO harmonics are altered as a function of $\overline{\Gamma}$. Figure 4 shows how the structure of the first three TO harmonics differ in a weak ($\overline{\Gamma}=10^{19}$ N m) versus a strong ($\overline{\Gamma}=10^{22}$ N m) gravitational coupling regime. The oscillation of the TC – entrained by the TO modes – is only weakly resisted by gravitational coupling when the latter is weak. In contrast, when gravitational coupling is strong, the oscillation of the TC is severely restricted. Using standing waves on a string as an analogy, a strong gravitational coupling changes the boundary condition on $\Omega_{f}$ from a free-end at $s\rightarrow 0$ to one more akin to a fixed-end at $s<s_{tc}$. Because the latter implies a shorter mode wavelength, the period of the modes is shorter, as seen in Figure 3a.
Figure 3a also shows that the point of transition of every TO mode between the weak and strong gravitational coupling regime is set by the period of the MICG mode. In other words, the TC remains gravitationally quasi-locked to the mantle for a given TO mode if its period is much longer than the MICG period. Hence, even though the MICG mode is no longer an individual mode of the system, its period retains an influence in the characteristics of the resulting TO modes. When the period of a TO mode approaches the MICG period, a resonance effect occurs, leading to a higher $Q$ (Figure 3b). The degree of resonant amplification depends on the attenuation of the TO modes caused by EM damping at the CMB but also by the viscous relaxation time of the inner core $\tau$. When the latter is of the same order or shorter than the period of a given TO mode, $Q$ is reduced. This is shown in Figure 3b, where we have computed $Q$ for $\tau=100$, $10$ and $3$ years.
@@
Figure 4: The angular velocity as a function of cylindrical radius ($s/r_{f}$) of the (a) first, (b) second, and (c) third TO harmonics, for $\overline{\Gamma}=10^{19}$ N m (blue lines for $\Omega_{f}$, thick light blue lines for $\Omega_{i}$) and $\overline{\Gamma}=10^{22}$ N m (red lines for $\Omega_{f}$, thick orange lines for $\Omega_{i}$). In each case, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
@@ 5 Discussion and conclusions
As our results show, when the magnetic field strength in the bulk of the outer core is of the order of a few mT as predicted from dynamo models (christensen06; aubert09), the TC is strongly coupled by EM stress to the fluid cylinders in the ROTC. Motion in the latter take the form of Alfvén waves. If Alfvén waves launched at the TC can traverse through the entire width of the ROTC before being attenuated, the MICG mode is no longer a distinct independent mode of the coupled Earth system. Instead, it gets absorbed into the spectrum of normal modes of TO. The MICG period retains a meaning, acting as a point of resonance for TO modes, and marking the transition from a TO mode in which the oscillating motion of the TC is weakly impeded by gravitational coupling to one where the TC remains quasi-locked to the mantle.
Though presented differently, our results are in agreement with those of mound03. In their study, which uses a broadly similar model as ours, they do not compute the modes of oscillations directly. Rather, they impose a forcing on the inner core, and monitor the response of the mantle angular velocity; frequencies at which the motion of the latter is amplified mark the location of normal modes. With a uniform $B_{s}$-field of 0.4 mT (similar to our case $c$), and with no EM coupling at the CMB, they find a set of resonances which are amplified in the broad vicinity of the MICG frequency associated with TO modes (see their Figure 4). This is in agreement with our results: we also find that TO modes have a higher $Q$ when their periods is in the vicinity of the MICG period. mound03 interpreted these results to imply that the MICG mode is complicated by its interaction with TO modes. By directly computing the modes of the coupled system, we have refined this interpretation; the MICG mode is no longer a distinct independent mode, it gets integrated into the spectrum of TO modes when Alfvén waves are weakly attenuated.
With a $B_{s}$ strength in the core expected to be of the order of a few mT, Alfvén waves can traverse the core in only a few years (gillet10). For the MICG mode to remain an independent mode of the coupled Earth system, the attenuation time of Alfvén waves by coupling at the CMB (EM or otherwise) would then need to be of the order of one year or less. As an indication, such a fast attenuation would require a mantle conductance larger than $10^{9}$ S, so large in fact that the gravest TO modes and the MICG mode would all have a $Q$ below 1.
This conclusion is important regarding the interpretation of the 5.9 year oscillation in the LOD, which includes a clear contribution from core (e.g. pfeffer23; cazenave25). A number of studies have suggested that this signal may represent the signature of the MICG mode (mound06; chao17; ding18; duan18; shih21). Our results show that this interpretation, which involves the ROTC playing a passive role, cannot be correct when the magnetic field strength inside the core is a few mT. Consequently, a measure of the gravitational coupling constant $\overline{\Gamma}$ cannot be inferred on the basis of the 5.9 year LOD signal, and in turn, it cannot be used as an indirect constraint on the amplitude of the degree 2 order 2 mantle mass anomalies. Interannual LOD changes involving the core must then be the manifestation of either TO modes, or any other free or driven core flow motion. This is consistent with the results of gillet10 who showed that, for the 5.9 year LOD signal, the mantle exchanges its angular momentum with the fluid core, and dominantly the ROTC given that latter has a much larger moment of inertia than the TC.
@@
The most recent estimates of the gravitational coupling constant $\overline{\Gamma}$ based on mantle flow models suggest a range of $0.3-2\times 10^{20}$ N m (davies14), corresponding to a range of MICG periods of 7 to 19 years. Assuming a $B_{s}$-field of approximately 3 mT or larger, all TO modes should then have a period shorter than the MICG period. This implies that TO modes should be weakly influenced by gravitational coupling in the sense that oscillations of the TC region should not be severely restricted by the mantle. Furthermore, the period of the TO modes, especially those of the first few harmonics, may be sufficiently close to the MICG period that they may be resonantly amplified. We thus expect that, at subdecadal periods, the TC (including the inner core) can undergo oscillating motion with respect to the mantle at a similar angular velocity amplitude as cylinders outside the TC, whether driven by TO modes or by any other core flow. At a period of 6 years, and based on the observed fluctuations of the magnetic field, zonal flow fluctuations inside the TC have an amplitude of approximately $v_{\phi}\sim 0.4$ km yr-1 (gillet10). Using this as a representative measure, this corresponds to longitudinal angular fluctuations of the inner core of approximately $0.02^{\circ}$ (lecomte23).
@@
Our results also have important implications for the interpretation of the subdecadal fluctuations in inner core rotations recently inferred by seismic observations, notably in the study of wang22 which suggests a super-rotation of the inner core of $\sim 0.29^{\circ}$ between 1971 and 1974 and a sub-rotation of $\sim 0.1^{\circ}$ between 1969 and 1971. wang22 further showed that the amplitude and phase of these inferred inner core rotation changes ($\Omega_{i}$) are consistent with those reconstructed from the 5.9 year LOD signal under the assumption that the latter is caused by the MICG mode. We first note that the inferred amplitude of inner core oscillations, if part of the MICG mode, would then imply similar fluctuations in zonal core flows inside the TC, and these are much larger than the approximate upper bound of $0.02^{\circ}$ stated above. It is possible that zonal core flows are not sufficiently well resolved inside the TC, so their amplitude may be larger. Alternatively, a part of the seismic signal could be explained by changes in the ICB topography (wen06; yao19), including from viscous relaxation (vidale25). In any case, its large amplitude notwithstanding, our results indicate that subdecadal changes in $\Omega_{i}$ cannot be associated with the MICG mode, contrary to the interpretation made by wang22. Instead, we favour an interpretation of subdecadal $\Omega_{i}$ changes entrained by TO modes or, more generally, by Alfvén waves. In this case, unlike for the MICG mode, a prediction of $\Omega_{i}$ cannot be constructed directly from the LOD signal but instead depends on many factors, including the relative strengths of the EM and gravitational torque on the mantle and the structure of Alfvén waves. It is of course possible that the phase of $\Omega_{i}$ entrained by Alfvén waves is similar to that predicted from the LOD signal on the basis of the MICG mode.
Finally, we have noted that the MICG period can act as a resonance to amplify TO modes. More generally any driven motion of the TC at a period that approaches that of the MICG would be amplified. If $\overline{\Gamma}\approx 3\times 10^{19}$ N m, the MICG period is approximately 20 years. If the inner core viscous relaxation time is of the same order or longer than this period, some of the observed multidecadal LOD changes caused by the core (e.g. finlay23) may then result from an amplification of driven zonal core flow motion (e.g. more18) due to their close temporal proximity to the MICG period.
@@ Acknowledgements.
This work was financially supported by a Discovery Grant from NSERC/CRSNG (Canada). Figures were produced using the GMT software (gmt).
@@
The codes necessary to reproduce all of our numerical experiments are freely accessible at the following data repository: dumberry25data.

### v2

@@ 4 Results
Figure 2: The angular velocity of the inner core $\Omega_{i}$ (thick light blue), mantle $\Omega_{m}\times 100$ (thick orange) and of the fluid core $\Omega_{f}$ (thin dark blue) as a function of cylindrical radius ($s/r_{f}$) for the modes identified by a purple outline on the respective panels of Figure 1. The dashed black line corresponds to the prediction $\Omega_{m}=-(C_{tc}/C_{m})\Omega_{i}$ (multiplied by 100) when the exchange of angular momentum is between the mantle and whole of TC. In each panel, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
The above analysis shows that the MICG only remains an independent mode when Alfvén waves launched by the oscillating TC get attenuated before they reach $s=r_{f}$ (either because $Lu<1$ or $Lu_{em}<1$). If so, they do not have a feedback on the motion of the TC, and the MICG mode can exist independently from the TO modes. In contrast, when Alfvén waves can travel through the entire width of the ROTC, the motion of the TC must be compatible with the structure of TO modes, and the MICG mode gets absorbed in the spectrum of TO modes. Alfvén waves may be partially or fully absorbed once they reach the equator of the CMB [44, 24], but the key factor is whether they can reach this point.
The $B_{s}$-field chosen for cases $b$ and $c$ above remains weak (smaller than 0.3 mT). The TO harmonics at periods close to the MICG mode have short wavelengths and the net angular momentum of the ROTC for these modes remains small. The angular momentum budget for these TO modes remains primarily an exchange between the mantle and the whole of the TC via gravitational coupling. Figures 2b and 2c show that $\Omega_{m}$ indeed remains approximately equal to $(C_{tc}/C_{m})\Omega_{i}$ for these selected modes. Hence, these TO modes share the same angular momentum dynamics as the MICG mode.
Although presented differently, our results are in agreement with those of [37]. In their study, which uses a similar model, they do not compute modes directly. Rather, they impose a forcing on the inner core, and monitor the response of the mantle. The frequencies at which the angular motion of the mantle is amplified mark the periods of the normal modes. With a uniform $B_{s}$-field of 0.4 mT (similar to our case $c$), and with no EM coupling at the CMB, they find (as we do) a set of TO modes amplified in the broad vicinity of the MICG frequency (see their Figure 4). This amplification leaves a broad peak in the LOD at the period of the MICG mode [39].
We now explore how the MICG mode affects the TO modes in a strong $B_{s}$-field regime. We restore $G_{m}=10^{8}$ S and set $\overline{B}_{s}=3.3$ mT, typical of the field strength expected inside the core [25]. With these choices, the fundamental TO mode has a period of approximately 6 years. Alfvén waves can readily propagate across the entire ROTC in a short timescale of $\tau_{A}=2.5$ yr. This is much faster than $\tau_{em}$ and also much faster than $\tau_{d}$ which (for a period of 6 yr) is now close to $3\times 10^{4}$ yr. Hence, both $Lu\gg 1$ and $Lu_{em}\gg 1$, and, accordingly, the spectrum of modes should not include an independent MICG mode; we have verified that this is indeed the case for any choices of $\overline{\Gamma}$. With our chosen mantle conductance, $Lu_{m}=0.34$; this is sufficiently high that Alfvén waves are partly absorbed at the equator of the CMB, yet low enough that distinct spectral peaks at the periods of the TO modes would remain visible in the LOD spectrum, compatible with observations [24].
@@
Figure 4: The angular velocity as a function of cylindrical radius ($s/r_{f}$) of the (a, d) first, (b, e) second, and (c, f) third TO harmonics for $\overline{\Gamma}=10^{19}$ N m (left column) and $\overline{\Gamma}=10^{22}$ N m (right column). Thin blue lines: $\Omega_{f}$, thick light blue lines: $\Omega_{i}$ and thick orange lines: $\Omega_{m}\times 100$. The dashed black line corresponds to the prediction $\Omega_{m}=-(C_{tc}/C_{m})\Omega_{i}$ (multiplied by 100). The dashed red line corresponds to the prediction of $\Omega_{m}=-L_{rotc}/C_{m}$ (multiplied by 100), where $L_{rotc}$ is the net angular momentum of the ROTC. In each panel, the phase is chosen when $\Omega_{i}$ is maximum in the prograde direction and the amplitude is normalized such that $\max|\Omega_{f}|=1$.
Figure 3 shows how the periods (panel a) and $Q$ (panel b) of the first four TO harmonics are altered as a function of $\overline{\Gamma}$. Figure 4 shows how the structure of the first three TO harmonics differ in a weak ($\overline{\Gamma}=10^{19}$ N m) versus a strong ($\overline{\Gamma}=10^{22}$ N m) gravitational coupling regime. When gravitational coupling is weak ($\overline{\Gamma}=10^{19}$ N m), the oscillation of the TC – entrained by the TO modes – is only weakly resisted by gravitational coupling. We also show on Figure 4 the predictions of $\Omega_{m}$ based on an angular momentum exchange with the TC ( $\Omega_{m}=-(C_{tc}/C_{m})\Omega_{i}$) and with the ROTC ($\Omega_{m}=-L_{rotc}/C_{m}$, where $L_{rotc}$ is the net angular momentum of the ROTC). The sum of these predictions matches $\Omega_{m}$ well, reflecting that the core-mantle angular momentum is conserved. But neither provides a good fit to $\Omega_{m}$, which has a much smaller amplitude. This indicates that the angular momentum budget for these modes is primarily an exchange between the TC and the ROTC; $\Omega_{m}$ reflects the small imbalance in the angular momenta of the TC and the ROTC. When gravitational coupling is strong ($\overline{\Gamma}=10^{22}$ N m), the oscillation of the TC is more restricted. Using standing waves on a string as an analogy, a strong gravitational coupling changes the boundary condition on $\Omega_{f}$ from a free-end at $s\rightarrow 0$ to one more akin to a fixed-end at $s<s_{tc}$. Because the latter implies a shorter mode wavelength, the period of the TO modes are shorter (see Figure 3a). The angular momentum budget is now dominantly an exchange between the mantle and the ROTC (although the TC plays a non-negligible role). Whether $\overline{\Gamma}$ is weak or strong, these TO modes have a different angular momentum dynamics than the MICG mode. Instead, the ROTC is involved in the angular momentum budget and it is the whole of the core that exchanges its angular momentum with the mantle.
Figure 3a also shows that the point of transition between the weak and strong gravitational coupling regime for every TO mode is set by the period of the MICG mode. Hence, even though the MICG mode is no longer an individual mode of the system, its period retains an influence in the characteristics of the resulting TO modes. When the period of a TO mode approaches the MICG period, a resonance effect occurs, leading to a higher $Q$ (Figure 3b). Note that the level of amplification is larger when the MICG period matches higher TO harmonics, consistent with results obtained in the weak $B_{s}$-field regime. When the MICG period is longer than the period of the fundamental TO harmonic, the level of amplification is limited. The amplification also depends on the attenuation of the TO modes caused by EM damping at the CMB and by the viscous relaxation time of the inner core $\tau$. When the latter is of the same order or shorter than the period of a given TO mode, $Q$ is significantly reduced. This is shown in Figure 3b, where we have computed $Q$ for $\tau=100$, $10$ and $3$ years.
@@ 5 Discussion and conclusions
Our results show that if Alfvén waves launched at the TC can travel across the entire width of the ROTC before being attenuated, the MICG mode is no longer a distinct independent mode of the core-mantle system. For a magnetic field strength in the bulk of the outer core equal to a few mT, as predicted by dynamo models [14, 3], Alfvén waves can readily traverse the core in only a few years [25], much shorter than their attenuation time. The MICG mode should be absorbed into the spectrum of TO modes. The MICG period retains a dynamical influence, as it acts as a point of resonance for TO modes, and marks the transition from a TO mode in which the oscillating motion of the TC is weakly affected by gravitational coupling to one in which the motion of the TC is strongly impeded by it.
For the MICG mode to remain an independent mode of the core-mantle system, the attenuation time of Alfvén waves by coupling at the CMB (EM or otherwise) would need to be of the order of one year or less. As an indication, such a fast attenuation requires a mantle conductance larger than $10^{9}$ S, so large in fact that the gravest TO modes and the MICG mode would all have a $Q$ below 1. Hence, the very fact that TO modes are excited to an observable level points to a limit in the mantle conductance which is well below $10^{9}$ S [24]. In addition, such a large conductance would introduce a significant delay between the time a magnetic signal is generated at the CMB and its observation at the surface [33, e.g.], yet no evidence of such a delay is observed [30, 26, e.g.]. The latest results from mineral physics and seismology also point to an upper bound on the mantle conductance well below $10^{9}$ S. The electrical conductivity of FeO at CMB conditions is in the range of $10^{4}-10^{5}$ S m-1 [41, 29], and while a thin layer of approximately 2 km of such iron-enriched material is supported by normal modes [43], a thicker layer of a few tens of km – such as would be required for a conductance $10^{9}$ S – is not. Given the limits on the possible mantle conductance, it is difficult to escape the conclusion that the MICG mode should be absorbed in the spectrum of TO modes.
This conclusion is important regarding the interpretation of the 5.9 year oscillation in the LOD, which includes a clear contribution from the core [42, 11, e.g.]. A number of studies have suggested that this signal may represent the signature of the MICG mode [39, 12, 17, 18, 45]. This interpretation rests on the assumption that the magnetic field in the core is weak (equal to a fraction of a mT). In such a case, the resonant excitation of a cluster of high TO harmonics with periods close to that of the MICG produces a distinct LOD peak at the MICG period [37, 39]. The angular momentum budget for these TO modes is primarily an exchange between the mantle and the whole of the TC, just as it is for the MICG mode. However, when the magnetic field strength inside the core is a few mT (as we expect), the subdecadal period range is now populated by the first few TO harmonics. These have distinct, separated periods and their angular momentum dynamics feature a leading order contribution from the ROTC; they do not behave like the MICG mode. These TO modes have a higher $Q$ if their periods are close to the MICG period, but the peaks in the LOD that they may leave correspond to the distinct periods of the TO modes, not the period of the MICG. Hence, for a strong magnetic field inside the core of a few mT, the 5.9 year LOD signal cannot represent a direct signature of the MICG mode. Instead, it must be a signature of the fundamental TO mode, or any other free or driven core flow motion. This is consistent with the results of [25] who showed that, for the 5.9 year LOD signal, the mantle exchanges its angular momentum with the whole of the core, not the TC alone.
@@
The most recent estimates of the gravitational coupling constant $\overline{\Gamma}$ based on mantle flow models suggest a range of $0.3-2\times 10^{20}$ N m [15], corresponding to a range of MICG periods between 7 and 19 years. Assuming a $B_{s}$-field of approximately 3 mT or larger, all TO modes should then have a period shorter than the MICG period. This implies that TO modes should be weakly influenced by gravitational coupling in the sense that oscillations of the TC region should not be severely restricted by the mantle. We thus expect that, at subdecadal periods, the TC (including the inner core) can undergo oscillating motion with respect to the mantle at a similar angular velocity amplitude as cylinders outside the TC. At a period of 6 years, and based on the observed fluctuations of the magnetic field, zonal flow fluctuations inside the TC have an amplitude of approximately $v_{\phi}\sim 0.4$ km yr-1 [25]. Using this as a representative measure, this corresponds to longitudinal angular fluctuations of the inner core of approximately $0.02^{\circ}$ [34].
@@
Our results also have important implications for the interpretation of the subdecadal fluctuations in inner core rotations recently inferred by seismic observations, notably in the study of [48] which suggests a super-rotation of the inner core of $\sim 0.29^{\circ}$ between 1971 and 1974 and a sub-rotation of $\sim 0.1^{\circ}$ between 1969 and 1971. [48] further showed that the amplitude and phase of these inferred inner core rotation changes ($\Omega_{i}$) are consistent with those reconstructed from the 5.9 year LOD signal under the assumption that the latter is caused by the MICG mode. The authors have themselves challenged their own interpretation in a subsequent study focused on the time window 1991-2023, in which a 6-yr inner core oscillation of a similar magnitude is not observed [49]. Based on our results, we favour an interpretation of subdecadal changes in $\Omega_{i}$ entrained by TO modes or, more generally, by Alfvén waves. In this case, unlike for the MICG mode, a prediction of $\Omega_{i}$ cannot be constructed directly from the LOD signal but instead depends on many factors, including the relative strengths of the EM and gravitational torque on the mantle and the structure of Alfvén waves. It is of course possible that the phase of $\Omega_{i}$ entrained by Alfvén waves is similar to that predicted from the LOD signal on the basis of the MICG mode. However, the amplitude of the subdecadal changes $\Omega_{i}$ should be limited to approximately $0.02^{\circ}$, as stated above, much smaller than the amplitudes inferred by [48].
Although our focus here has been on the MICG mode, our model may be adapted to monitor the angular momentum response of the core-mantle system to a specified forcing in the core, either at a given frequency [37, 38, e.g.] or for a stochastic forcing [24, e.g.]. The parameter space of $\overline{\Gamma}$, $\tau$ and $G_{m}$ may be surveyed to determine the range of values that allow to best reproduce the characteristics of the observed Alfvén waves, while also matching the amplitude and phase of the LOD changes at subdecadal and decadal timescales. This would provide constraints on these parameters, which in turn would provide further constraints on the structure, composition and dynamics of the lower mantle and inner core.
@@ Acknowledgements.
I wish to thank Jon Mound and an anonymous reviewer for comments and suggestions that have greatly improved the quality of this paper. This work was financially supported by a Discovery Grant from NSERC/CRSNG (Canada). Figures were produced using the GMT software [50].
@@
Codes to reproduce all numerical experiments are freely accessible at the following data repository: [19].